The line x-1/3=y+1/4=z-2/3 passes through the point (4,2,3).


1
Expert's answer
2021-04-13T17:32:12-0400

if line "\frac{x-1}{3} =\frac{y+1}{4} = \frac{z-2}{3}" passes trough the point (4,2,3) then "\frac{4-1}{3} =\frac{2+1}{4} = \frac{3-2}{3}" , but

1 "\ne" 0.75"\ne" "\frac{1}{3}" . so the line doesn't passes through the point(4,2,3)